11. Sum of the areas of two squares is 468 m2. If the difference of their perimeters is 24 m, find the sides of the two squares.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Let the sides of the two squares be x m and y m. Therefore, their perimeter will be 4x and 4y respectively And area of the squares will be x2 and y2 respectively. Given, 4x – 4y = 24 x – y = 6 x = y + 6 Also, x2 + y2 = 468 ⇒ (6 + y2) + y2 = 468 ⇒ 36 + y2 + 12y + y2 = 468 ⇒ 2y2 + 12y + 432 = 0 ⇒ y2 + 6y – 216 = 0 ⇒ y2 + 18y – 12y – 216 = 0 ⇒ y(y +18) -12(y + 18) = 0 ⇒ (y + 18)(y – 12) = 0 ⇒ y = -18, 12 As we know, the side of a square cannot be negative. Hence, the sides of the squares are 12 m and (12 + 6) m = 18 m.